A bank pays you 4.5% interest per year. In addition, you receive |100 as bonus at

the end of the year (after the interest is paid). Find a recurrence for the amount of

money after n years if you invest |2000.



1
Expert's answer
2021-04-15T06:31:31-0400

The amount of money after 1 year


"A_1=2000(1+0.045)+100"

Let "A_{n-1}=" the amount of money after "n-1" years, "A_n=" the amount of money after "n" years. Then


"A_n=A_{n-1}(1+0.045)+100, n\geq2"

A recurrence for the amount of money after n years 


"A_1=2000(1+0.045)+100=2190"

"A_n=A_{n-1}(1+0.045)+100, n\geq2"